OpenStudy (anonymous):

A store has a total of 15 bicycles and tricycles. There are 37 wheels. How many tricycles does the store have?

OpenStudy (anonymous):

plz answer

OpenStudy (anonymous):

need answer now

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

15 * 2 = 30 37 - 30 = 7 \[\frac{ 7 }{ 3 } = 2\frac{ 1 }{ 3 }\ Tricycles]

OpenStudy (anonymous):

sooo its 7??

OpenStudy (anonymous):

@tom99

OpenStudy (anonymous):

\[\frac{ 7 }{ 3 } = 2\frac{ 1 }{ 3 }\] tricycles

OpenStudy (anonymous):

doesn't add up :(

OpenStudy (anonymous):

12 is the answer

OpenStudy (anonymous):

oh... i though it said 15 bicycles and some tricycles

OpenStudy (anonymous):

I think